the difference of the digits of a 2 digit number is 2. the value of the number is 3 less than 5 times the sum of its digits
Answer by sudhanshu_kmr(1152)


let assume digits unit digit is y and tens digit is x.
case 1: when x-y = 2 (i.e x is greater than y)
10x+y = 5(x+y)-3
=> 10x + y = 5x+ 5y -3
=> 5x -4y = -3
after solving...
x = -11, y = -13
it is not possible...

case 2: when y -x =2 (i.e y is greater than x)

10x+y = 5(x+y) -3

after solving value of x and y.
x = 5, y = 7

so, required number is : 57